If a rectangular region having an area of 500 square units has its length increased by 20% and its width decreased by 10%, what is the area of the newly formed region?
Answer by stanbon(75887)

If a rectangular region having an area of 500 square units has its length increased by 20% and its width decreased by 10%, what is the area of the newly formed region?
----
Origin Rectangle::
L*W = 500
------
New Rectangle::
1.2*L * 0.9W = 1.08*L*W = 1.08*500 = 540 sq units
-----------------------
Cheers,
Stan H.
------------------